The math club and science club had fundraisers to buy supplies for a hospice, The math club spent $135 buying sixcases of juice and one case of bottled water. The science club spent $110 buying four cases of juice and two cases of bottled water. How much did a case of juice cost? How much did a case of bottled wated cost?
Answer by stanbon(75887)

The math club and science club had fundraisers to buy supplies for a hospice, The math club spent $135 buying sixcases of juice and one case of bottled water. The science club spent $110 buying four cases of juice and two cases of bottled water. How much did a case of juice cost? How much did a case of bottled wated cost?
------
Equations:
6j + w = 135
4j + 2w = 110
---------------------------
Modify for elimination:
12j + 2w = 270
12j + 6w = 330
-----
4w = 60
w = 15 (water costs $15 per bottle)
-----------------------------------------
Cheers,
Stan H.
---------------------------